### v1

@@ 2 Model
In Section 2.1 we present our contract setting; we reduce this setting to a purely combinatorial setting in Section 2.2.
2.1 Contract Design Setting
  • $v_{i}$: The reward which the principal obtains if agent $i$ completes their job before the deadline.
  • $\{(F_{i}^{j},c_{i}^{j})\}_{j\in[m]}$: The possible effort levels of agent ii. [^1] All agents have the same number of effort levels without loss of generality, since we can add “dummy” effort levels that would never be incentivized by the principal. Each effort level $j\in[m]$ determines a distribution $F_{i}^{j}$ over the time it takes the agent to complete the job, as well as a cost $c_{i}^{j}\geq 0$. If the agent chooses to exert effort level $j$, they incur cost $c_{i}^{j}$, and complete their job within time $s_{i}$ drawn (independently) from distribution $F_{i}^{j}$. We assume non-decreasing costs $c_{i}^{1}\leq\dots\leq c_{i}^{m}$. Let $\mathbf{F}_{i}=\left(F_{i}^{1},\ldots,F_{i}^{m}\right)$. Let $\mathrm{supp}(F_{i}^{j})$ denote the support of $F_{i}^{j}$, and $\mathrm{supp}(\mathbf{F}_{i})=\cup_{j\in[m]}\mathrm{supp}(F_{i}^{j}).$ We assume the support of each distribution is finite and is explicitly given.
Definition 2.1 (Implementable effort level).
$\min\quad$ $\sum_{s\in\mathrm{supp}(F_{i}^{j})}t_{i}(s)\cdot F_{i}^{j}(s)$
$t_{i}(s)\geq 0$ $\forall s\in\mathrm{supp}(\mathbf{F}_{i})$
The principal’s problem can then be described as: when hiring agent $i$, the principal chooses an effort level $j$ to incentivize, and offers the optimal contract that incentivizes that effort level. The principal’s collected reward is the sum of the rewards of the jobs that the agents get to complete with a total processing time less than the budget $S$. The principal aims to maximize their profit: the expected total reward minus the expected sum of transfers made to the agents.
Adaptive versus non-adaptive policies.
@@ 2.1 Contract Design Setting
(Fully-Adaptive Policies). A fully-adaptive policy, $\varphi:2^{I}\times[0,1]\rightarrow[n]\times[m]\cup\{\perp\}$, takes as input the unassigned jobs $J\subseteq I$ and the remaining time budget $S^{\prime}\leq 1$, and chooses which agent to hire next and at which implementable effort level. This procedure is repeated until:
  • an agent finishes their job after the deadline; or
  • the policy decides not to execute another job (outputs $\perp$ as the next job).
Let $\mathrm{profit}(\varphi)$ be a random variable representing the profit obtained by the principal when using policy $\varphi$ on instance $I$. We denote by $\mathrm{ADAPT}(I)=\max_{\varphi}\operatorname{\mathbb{E}}[\mathrm{profit}(\varphi)]$ the optimum expected profit obtained by an adaptive policy for instance $I$. When referring to an algorithm $\mathcal{A}$ that is given an instance $I$ and returns a policy $\varphi_{I}$, we denote by $\mathcal{A}(I)=\mathrm{profit}(\varphi_{I})$ the random variable representing the profit obtained by the principal when using the policy outputted by algorithm $\mathcal{A}$’s policy on instance $I$.
@@
(Non-Adaptive Policies) A non-adaptive policy is an ordering of agents $\pi:[n]\rightarrow[n]\cup\{\perp\}\times[m]$. For some $i\in[n]$, $\pi(i)$ outputs the $i$th selected agent and the effort level the policy incentivizes the agent to exert.
Let $\Pi$ be the set of non-adaptive policies. We denote by $\mathrm{NON\_ADAPT}(I)=\max_{\pi\in\Pi}\operatorname{\mathbb{E}}[\mathrm{profit}(\pi)]$ the optimum expected profit obtained by a nonadaptive policy for instance $I$.
@@
(Stopping-Time Adaptive Policies) Similarly to non-adaptive policies, a stopping-time adaptive policy chooses a non-adaptive ordering of the agents and their effort levels $\pi$. However, it may adaptively choose to stop at any point. Without loss of generality, we assume the policy stops if every remaining sequence of agents yields a negative expected profit.
Let $\Pi^{\prime}$ be the set of stopping-time adaptive policies. $\mathrm{STOP\_ADAPT}(I)=\max_{\pi\in\Pi^{\prime}}\operatorname{\mathbb{E}}[\mathrm{profit}(\pi)]$ is the optimum expected profit obtained by a stopping-time policy for instance $I$.
2.2 Multi-Choice Stochastic Knapsack Problem With Costs
We define a generalization of Stochastic Knapsack, where items have multiple choices of assignment, and each is associated with a size distribution, and a cost. We show that our Knapsack Contracts problem reduces to this problem. This implies that our algorithmic results can be obtained in this setting, and the results will carry through to the contracts setting.
Definition 2.5 (Multi-Choice Stochastic Knapsack Problem With Costs).
An instance of the multi-choice stochastic knapsack problem with costs (or MSKC for short) consists of $n$ items, where each item has $m$ choices, each choice has a different cost and size distribution. That is $I=\{v_{i},\{p_{i}^{j},F_{i}^{j}\}_{j\in[m]}\}_{i\in[n]}.$ An item $i$ can be selected with a single choice $j$. In this case, the cost $p_{i}^{j}$ must be incurred up front, and the item’s size is realized from distribution $F_{i}^{j}$. There is a budget of size $S$ ($S=1$ without loss of generality). A policy selects item-choice pairs sequentially, and terminates either when it decides not to select any item, or when the total size of selected items exceeds the budget. The profit of a policy is the total value of items selected that do not exceed the budget minus the total cost of the selected items, including the cost of an item that exceeds the budget, if one exists.
Adaptive, Stopping-time Adaptive and Non-adaptive policies for the MSKC problem are defined analogously to the way they are defined for the Knapsack Contracts problem. The following lemma (Proof in Appendix F) shows that in order to find an approximately optimal policy for the Knapsack Contracts problem, it suffices to find an approximately optimal policy for the MSKC problem.
Lemma 2.1 (Reduction to Multi-Choice Stochastic Knapsack with Costs).
Given an efficient algorithm that finds a policy that $c$-approximates the optimal policy MSKC problem, one can efficiently find a policy that $c$-approximates the optimal policy for the knapsack contracts problem.
2.3 Inverse of Return on Investment Parameter
Consider an instance of the Knapsack Contracts problem. For a job and implementable effort level pair, $(i,j)$, let
$\tilde{t}_{i}^{j}=\operatorname{\mathbb{E}}_{s\sim F_{i}^{j}}[t_{i}(s)],$
where $t_{i}$ is the contract that minimizes the expected payment while incentivizing effort level $j$, as given by the LP in Definition 2.1. That is, $\tilde{t}_{i}^{j}$ is the minimal expected transfer required in order to incentivize agent $i$ to take action $j$.
The hardness results presented in the paper critically exploit problem instances characterized by pairs (job, action) with high values $v_{i}$ and expected transfers $\tilde{t}_{i}^{j}$, but a comparatively low net profit
$v_{i}\cdot\Pr_{s_{i}^{j}\sim F_{i}^{j}}[s_{i}^{j}\leq 1]-\tilde{t}_{i}^{j}.$
In such cases, a job may yield a small positive profit of $v_{i}-\tilde{t}_{i}^{j}$ when the knapsack is empty, yet incur a large negative penalty of up to $\tilde{t}_{i}^{j}$ when capacity is low. Moreover, this fluctuation can be abrupt: even an $\epsilon$ increase in load may decrease the expected profit to $\tilde{t}_{i}^{j}$. We show that this drastic fluctuation property is an inherent obstacle in our model. To do so, we define the Inverse of the Return on Investment parameter:
Definition 2.6 (Inverse of ROI parameter).
We define
$\mathrm{ROI}(i,j)\ =\ \frac{v_{i}\cdot\Pr_{s_{i}^{j}\sim F_{i}^{j}}[\,s_{i}^{j}\leq 1\,]\;-\;\tilde{t}_{i}^{j}}{\tilde{t}_{i}^{j}},$
and
$\alpha\;=\;\max_{i,j\in\mathsf{I}_{i}}\frac{1}{\mathrm{ROI}(i,j)}\,$,
where $\mathsf{I}_{i}$ is the set of implementable actions for job $i$. We sometimes refer to $\alpha$ as the Inverse of ROI (or IOR) parameter.
Intuitively, a large value of $\alpha$ signifies the presence of at least one job where for some effort level, the transfer required to incentivize it is much larger than the net profit from successfully executing the job before the deadline. We show a $\Theta(\alpha)$ adaptivity gap and an $\Omega(\alpha)$ information gap (Section 5), suggesting that $\alpha$ captures precisely the “hardness” aspect for our problem.
Analogously, we define the IOR parameter for the MSKC problem as in Definition 2.6, but using $p_{i}^{j}$ instead of $\tilde{t}_{i}^{j}$. Given a Knapsack Contracts instance with an $\alpha$ IOR parameter, applying the reduction in Lemma 2.1 yields an MSKC instance with the same $\alpha$ as the analogous IOR parameter. This implies that any approximation as a function of $\alpha$ a policy in the MSKC instance would obtain will imply the same approximation as a function of $\alpha$ in the Stochastic Knapsack instance.
3 $O(\alpha)$-Approximate Non-Adaptive Contracts
In this section, we show how to find a non-adaptive set of contracts $\pi:[n]\to[n]\times[m]\cup\{\perp\}$ that obtains an $O(\alpha)$ approximation to the optimal profit the principal can obtain by any (possibly highly adaptive) policy.
By Lemma 2.1, we focus on finding an $O(\alpha)$-approximate non-adaptive policy for the MSKC problem, which yields an $O(\alpha)$-approximate non-adaptive set of contracts for the Knapsack Contracts problem. Our main theorem is the following:
@@ 3 $O(\alpha)$-Approximate Non-Adaptive Contracts
There exists a polynomial time algorithm for MSKC that computes for every input $I$ a non-adaptive policy $\pi_{I}:[n]\to[n]\times[m]\cup\{\perp\}$ such that
$\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathrm{profit}(\pi_{I})]}\leq O(\alpha).$
We first define an LP that both guides our solution and gives an upper bound on the performance of every adaptive policy. For every $i\in[n],j\in[m]$ and for a knapsack of size $t$, denote by $w_{ij}=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq t]-p_{i}^{j}$ and $\mu_{ij}=\operatorname{\mathbb{E}}_{s_{i}^{j}\sim F_{i}^{j}}[\min\{s_{i}^{j},t\}]$. The first parameter can be viewed as a proxy to the profit from job $i$ when executed at effort level $j$ and the second parameter is the expected truncated size.
@@
Note that every solution returned by the Simplex algorithm satisfies the above properties; the basic feasible solution of the LP has at most $n+1$ non-negative variables, and the constraints ensure that for each item $i$, there is at least one choice $j$ with $x_{ij}>0$, hence at most one item has two non-zero choices. However, the Simplex algorithm may run at exponential time in the worst case. Thus, we describe a polynomial time algorithm to compute such a solution.
First, for each item $i$, we characterize the choices $j$ for which there exists an optimal solution with $x_{ij}=0$. We then show that there always exists an optimal solution in which at most two choices have positive values.
@@
$(1)$ $\quad w_{ij^{\prime}}\geq w_{ij}$
$(2)$ $\quad\mu_{ij^{\prime}}\leq\mu_{ij}$
@@
First, if $j^{\prime}$ single-dominates $j$, define a new solution by setting
$x^{*}_{ij^{\prime}}=x_{ij^{\prime}}+x_{ij},\quad x^{*}_{ij}=0~.$
@@
This again yields a feasible solution whose objective value is at least as large as that of $x$ (see Fig. 1 for illustration, and Appendix E for the full proof). Therefore, in either case, there exists an optimal solution in which $x_{ij}=0$.
@@
There exists an optimal solution $x$ to $\Phi(t)$ in which, for every item $i$, at most two choices satisfy $x_{ij}>0$.
@@
Let $x$ be an optimal solution in which no choice is dominated; such a solution exists by Lemma 3.3. Without loss of generality we may further assume that no additional choices remain and that
$w_{i,j}\;<\;w_{i,j+1}\qquad\text{for every item }i\text{ and every }j\in\{1,\dots,m-1\}.$
Lemma 3.3 then also yields
$\mu_{i,j}\;<\;\mu_{i,j+1}\qquad\text{for all }i\text{ and }j\in\{1,\dots,m-1\}.$
@@
Suppose for some item $i$ there are strictly more than two choices that satisfy $x_{ij}>0$. Define $\tilde{\mu}=\sum_{j}x_{ij}\mu_{ij}$. Let $j^{\prime}=\arg\max_{j}\bigl{\{}\mu_{ij}\leq\tilde{\mu}\bigr{\}}$ and $j^{\prime\prime}=\arg\min_{j}\bigl{\{}\mu_{ij}\geq\tilde{\mu}\bigr{\}}$. By the definition of $\tilde{\mu}$ and by our assumption that at least two choices satisfy $x_{ij}>0$, such indices $j^{\prime}$ and $j^{\prime\prime}$ always exist. Set $\lambda=\frac{\mu_{ij^{\prime\prime}}-\tilde{\mu}}{\mu_{ij^{\prime\prime}}-\mu_{ij^{\prime}}}$. Consider the solution defined by $x^{*}_{ij^{\prime}}=\lambda$, $x^{*}_{ij^{\prime\prime}}=1-\lambda$, and $x^{*}_{ij}=0$ for all $j\notin\{j^{\prime}j^{\prime\prime}\}$. Note that this construction satisfies $\sum_{j}x^{*}_{ij}\,\mu_{ij}=\tilde{\mu}$. Therefore, it yields a feasible solution. Since the resulting point lies on the boundary of the convex set, its objective value is at least as large as that of the original solution $x$, (see Fig. 2 for illustration).
@@
By Lemmas 3.3, 3.4, it follows that there exists an optimal solution $x$ where at most two choices have positive values.
@@
Figure 3: For an optimal solution $x$, items $i_{1}$ and $i_{2}$ each have two choices with nonzero values. The solution for each item is a convex combination of the two choices, corresponding to the red dots in the graphs $\Gamma_{i_{1}}$ and $\Gamma_{i_{2}}$, which lie on the segments connecting the two choices. As the slope of the first item is larger, the modified solution indicated by the dashed arrows illustrates one possible update that maintain feasibility and increases the objective value.
For a set of item choices $\mathcal{S}\subset[n]\times[m]$, denote $size(\mathcal{S})=\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}$ and $\mu(S)=\sum_{(i,j)\in\mathcal{S}}\mu_{ij}$. The following Lemmas are direct extensions of key lemmas from Dean et al. [11], originally established for the classic Stochastic Knapsack problem with a budget of 1. We adapt these lemmas to our knapsack variant with multi-choice and costs, and for a general budget $t$. In this section, we only need the case where $t=1$; however, in Section 6 we use the generalized versions. The proofs are presented in Appendix E for completeness.
@@
For any set of item choices $\mathcal{S}$:
$\Pr[size(\mathcal{S})<t]\geq 1-\frac{1}{t}\mu(\mathcal{S}).$
@@
For any adaptive policy for the MSKC problem with budget $t$, let $\mathcal{S}$ be the random set of item choices that the policy inserts, then $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]\leq 2t$.
@@
For any constant $c\geq 2$ and budget $t$,
$\mathrm{ADAPT}\leq\Phi(ct)\leq c\cdot\Phi(t)~,$
where $\mathrm{ADAPT}$ is the optimal adaptive policy for budget $t$.
We are now ready to present our algorithm, SKC (Stochastic Knapsack with Costs):
@@
Solve linear program $\Phi(1)$ to obtain optimal solution $x$ with the structure from Lemma 3.2;
Let $\mathcal{C}=\{(i,j)\mid x_{ij}>0\}$ be the candidate item-choice pairs;
@@
   Remove $(i^{*},j)$ and $(i^{*},j^{\prime})$ from $\mathcal{C}$;
   Create virtual item $(i^{*},j^{*})$ with:
@@
   Add $(i^{*},j^{*})$ to $\mathcal{C}$;
 // Randomly select between original choices
   When attempting to insert $(i^{*},j^{*})$, choose $j$ with probability $x_{i^{*},j}$ and $j^{\prime}$ with probability $x_{i^{*},j^{\prime}}$;
end if

### v2

@@ 2 Model
In this section we introduce our problem, the Inverse ROI parameter, and our policy hierarchy.
2.1 The Multi-choice Stochastic Knapsack with Costs Problem (MSKC)
We generalize Stochastic Knapsack along two axes: (i) items can have multiple tiers, called choices, and (ii) each choice is associated with a cost. An instance $I=\{v_{i},\{p_{i}^{j},F_{i}^{j}\}_{j\in[m]}\}_{i\in[n]}$ of the multi-choice stochastic knapsack problem with costs (or MSKC for short), consists of $n$ items. Each item $i\in[n]$ has a value $v_{i}$ and $m$ choices. Each choice $j\in[m]$ has a (deterministic) cost $p_{i}^{j}$, and a size distribution $F_{i}^{j}$. There is also a knapsack capacity of size $S$, but without loss of generality, $S=1$.
When item $i$ is selected for the knapsack, a single choice $j$ must be specified. We refer to these as the item-choice pair (sometimes referring just to the item when the choice is clear from context). In this case, the cost $p_{i}^{j}$ is incurred up-front, and the item’s size $s_{i}$ is realized from distribution $F_{i}^{j}$. Item $i$’s realized utility is then $v_{i}-p_{i}^{j}$ (more precisely, this is the “potential” realized utility, i.e., the utility that will be gained if the item fits in the knapsack).
For a given instance $I$, a policy $\varphi$ possibly inspects the knapsack, then either selects an item-choice pair, or terminates. Item-choice pairs are selected according to the policy sequentially, until the total size of selected items exceeds the capacity, or until the policy decides to terminate. In the former case, we refer to the last selected item as the overflow item.
The realized utility of the policy is the total realized utility of items selected before the cumulative size exceeded the capacity, minus the cost of the overflow item, if one exists. Let $\mathrm{utility}(\varphi)$ be a random variable representing the utility obtained by the policy $\varphi$ on instance $I$. We also design algorithms that return policies. When referring to an algorithm $\mathcal{A}$ that is given an instance $I$ and returns a policy $\varphi_{I}$, we denote by $\mathcal{A}(I)=\mathrm{utility}(\varphi_{I})$ the random variable representing the utility obtained by the policy on instance $I$.
We are also interested in the expected utility of a pair $(i,j)$ of item and its choice. This depends on the remaining capacity $t\in[0,1]$, and is equal to $v_{i}\cdot\Pr_{s_{i}\sim F_{i}^{j}}[s_{i}\leq t]\;-\;p_{i}^{j}.$ Let
$w_{i,j}:=v_{i}\cdot\Pr_{s_{i}\sim F_{i}^{j}}[s_{i}\leq 1]-p_{i}^{j}$
be the expected utility of inserting $(i,j)$ when the knapsack is empty ($t=1$).
The lower bounds in this paper critically exploit instances with item-choice pairs $(i,j)$, which have high values and high costs, but only low expected utilities:
$0<w_{i,j}\ll\min\{v_{i},p_{i}^{j}\}.$
In such cases, an item-choice pair may yield a small positive expected utility when the knapsack is empty, yet incur a large negative expected utility, of up to $p_{i}^{j}$, when the capacity becomes limited. We show that this drastic instability property is an inherent obstacle in our model. To do so, we define the Inverse of the Return on Investment parameter:
Definition 2.1 (ROI and $\alpha$ Parameters).
We define
$\mathrm{ROI}(i,j)\ :=\ w_{i,j}/p_{i}^{j},~~~\alpha\ :=\ \max_{i,j}\{1/\mathrm{ROI}(i,j)\}.$
We also refer to $\alpha$ as the Inverse of ROI (IOR) parameter.
Intuitively, a large value of $\alpha$ signifies the presence of at least one item-choice pair where the upfront cost of the item greatly exceeds the utility from successfully inserting the item. Our results show that $\alpha$ is the key parameter governing the adaptivity gaps between different levels of adaptivity, which we now specify.
2.2 Adaptiveness Hierarchy
For a fixed instance, recall that a policy is applied iteratively, possibly observing the state of the knapsack in each iteration, in order to sequentially fill the knapsack. After the formal definitions we give an informal interpretation.
@@ 2.2 Adaptiveness Hierarchy
(Non-Adaptive Policy) A non-adaptive policy $\varphi:[n]\rightarrow[n]\times[m]\cup\{\perp\}$ takes as input the iteration number, and returns an item-choice pair, or $\perp$ for termination. Let $\mathrm{NON\_ADAPT}(I)$ be the optimum expected utility obtained by a non-adaptive policy for instance $I$.
@@
(Stopping-Adaptive Policy) A stopping-adaptive policy $\pi:[n]\times[0,1]\rightarrow[n]\times[m]\cup\{\perp\}$ takes as input the iteration number and the remaining knapsack capacity, and returns an item-choice pair, or $\perp$ for termination.
Let $\mathrm{STOP\_ADAPT}(I)$ be the optimum expected utility obtained by a stopping-adaptive policy for instance $I$.
@@
(Skipping-Adaptive Policy) A skipping-adaptive policy $\pi:[n]\times[n]\times[0,1]\rightarrow[n]\times[m]\cup\{\perp\}$ takes as input the iteration number, the last item in the knapsack and the remaining knapsack capacity, and returns an item-choice pair, or $\perp$ for termination. Let $\mathrm{SKIP\_ADAPT}(I)$ be the optimum expected utility obtained by a skipping-adaptive policy for instance $I$.
Definition 2.5.
(Fully-Adaptive Policy). A fully-adaptive policy $\varphi:[n]\times 2^{[n]}\times[0,1]\rightarrow[n]\times[m]\cup\{\perp\}$ takes as input the iteration number, the item set in the knapsack and the remaining knapsack capacity, and returns an item-choice pair, or $\perp$ for termination. Let $\mathrm{ADAPT}(I):=\max_{\varphi}\operatorname{\mathbb{E}}[\mathrm{utility}(\varphi(I))]$ be the optimum expected utility obtained by a fully-adaptive policy for instance $I$.
Intuitively, a non-adaptive policy is an ordering of items. For each $i\in[n]$, it specifies the $i$-th selected item and its associated choice, or $\perp$ for termination. A stopping-adaptive policy, like a non-adaptive one, is an ordering of items, but it can adaptively stop at any point. A skipping-adaptive policy is like a stopping-adaptive policy, with the additional ability to skip an item before stopping. It, too, is an ordering of items, and given the last inserted item it scans the successive ones in the ordering and chooses the next one to insert. [^1] Related to skipping-adaptive policies, we can consider an ordering of items and target capacities — the maximum remaining capacity with which each item will be entered into the knapsack. When given the last item inserted, the policy finds the subsequent item in the ordering, whose target capacity is below the remaining knapsack capacity. [10] study fully-adaptive (and non-adaptive) policies.
We remark that each policy can be represented by a decision tree; a fully-adaptive policy needs an exponentially-sized tree, but the other policies require only poly-sized trees. In fact stopping-adaptive and non-adaptive policies correspond to linear-sized decision trees.
3 Tight $\Theta(\alpha)$ Adaptivity Gap
In this section we establish a tight $\Theta(\alpha)$ gap between simple non-adaptive policies and the optimal fully-adaptive policy. Theorem 3.1 constructs an instance with an $\Omega(\alpha)$ gap between the best non-adaptive policy and the best stopping-adaptive policy. Theorem 3.2 gives a polynomial-time algorithm that outputs a non-adaptive policy whose expected utility is within an $O(\alpha)$ factor of the optimal fully-adaptive utility.
In Appendix F, we study an even more demanding setup, where the policy chooses an ordered set of jobs in advance, and incurs the cost of all of them, not only jobs that are inserted before the capacity is met. In the setting without costs (e.g., [10]), this setup is equivalent to a non-adaptive policy, as all costs are 0. In the MSKC Problem, we show a separation between the two. In Appendix F we define non-stopping policies, and show that there is an instance with an $\Omega(\alpha)$ gap between the optimal non-stopping policy and the optimal non-adaptive policy, and an $O(\alpha)$ approximation non-stopping algorithm to the optimal fully-adaptive policy.
3.1 Lower Bound
We show that there exists an $\Omega(\alpha)$ gap between the optimal Non-Adaptive policy and the optimal Stopping-Adaptive policy. Recall that both models require the algorithm to choose an item ordering at the beginning; the stopping-adaptive policy can then adaptively choose when to halt, while the non-adaptive policy must attempt all items in the chosen order, until all items are inserted or an overflow occurs.
@@ 3.1 Lower Bound
For every $\alpha$ there exists an instance $I$ of the MSKC problem such that $\frac{\mathrm{STOP\_ADAPT}(I)}{\mathrm{NON\_ADAPT}(I)}=\Omega(\alpha).$
We give a proof sketch here and defer the full proof of theorem 3.1 to section D.1.
Proof sketch.
Fix parameters $\varepsilon,\gamma\in(0,\tfrac{1}{2})$ and let $n=\lceil 1/\varepsilon\rceil$. Consider capacity $1$ and $n$ items with value $v=2$ and cost $p=2-\varepsilon$, so each successful insertion yields net gain $w=v-p=\varepsilon$ and hence $\alpha=\max_{i}p/w=(2-\varepsilon)/\varepsilon=\Theta(1/\varepsilon)$. For item $i\in[n]$, define a two-point size distribution
$s_{i}=\begin{cases}a_{i}&\text{with prob. }1-\varepsilon,\\ b_{i}&\text{with prob. }\varepsilon,\end{cases}\qquad a_{i}=\gamma^{\,n-i+1},\qquad b_{i}=1-\sum_{j<i}a_{j}.$
Since $\gamma<\tfrac{1}{2}$, the sequence $(a_{i})$ grows geometrically and satisfies $a_{i^{\prime}}>\sum_{j<i}a_{j}$ for all $i<i^{\prime}$, implying $b_{i}+a_{i^{\prime}}>1$. Thus, once any earlier item realizes its “big” size $b_{i}$, every subsequent attempt overflows even if it realizes “small”. A stopping-adaptive policy that attempts items in order and halts after the first big realization inserts a geometric number $X$ of items (truncated at $n$) with
$\mathbb{E}[X]=\sum_{t=1}^{n}(1-\varepsilon)^{t-1}=\frac{1-(1-\varepsilon)^{n}}{\varepsilon}\geq\frac{1-e^{-1}}{\varepsilon},$
and therefore achieves expected utility at least $\mathbb{E}[X]\cdot\varepsilon\geq 1-e^{-1}=\Omega(1)$.
In contrast, we show (in the full proof) that the expected utility of any non-adaptive policy that attempts inserting at least $k>1$ items has negative utility. Hence the optimal non-adaptive policy attempts only one item and attains $\mathrm{NON\_ADAPT}=\varepsilon$. Consequently,
$\frac{\mathrm{STOP\_ADAPT}}{\mathrm{NON\_ADAPT}}\geq\frac{1-e^{-1}}{\varepsilon}=\Omega(\alpha),$
as claimed. ∎
3.2 Upper Bound
In this section, we present an $O(\alpha)$-approximate non-adaptive policy for the MSKC problem. Our main theorem is the following:
Theorem 3.2.
There exists a polynomial time algorithm for MSKC that computes for every input $I$ a non-adaptive policy $\pi_{I}:[n]\to[n]\times[m]\cup\{\perp\}$ such that $\frac{\mathrm{ADAPT}(I)}{\operatorname{\mathbb{E}}[\mathrm{utility}(\pi_{I})]}\leq O(\alpha).$
We first define an LP that both guides our solution and gives an upper bound on the performance of every adaptive policy. For every $i\in[n],j\in[m]$ and for a knapsack of size $t$, denote by $w_{ij}=v_{i}\cdot\Pr_{{s_{i}^{j}\sim F_{i}^{j}}}[s_{i}^{j}\leq t]-p_{i}^{j}$ and $\mu_{ij}=\operatorname{\mathbb{E}}_{s_{i}^{j}\sim F_{i}^{j}}[\min\{s_{i}^{j},t\}]$. The first parameter can be viewed as a proxy to the utility from item $i$ inserted with choice $j$ and the second parameter is the expected truncated size.
@@ 3.2 Upper Bound
Note that every solution returned by the Simplex algorithm satisfies the above properties; the basic feasible solution of the LP has at most $n+1$ non-negative variables, and the constraints ensure that for each item $i$, there is at least one choice $j$ with $x_{ij}>0$, hence at most one item has two non-zero choices. However, the Simplex algorithm may run in exponential time in the worst case. Thus, we describe a polynomial time algorithm to compute such a solution.
First, for each item $i$, we characterize the choices $j$ for which there exists an optimal solution with $x_{ij}=0$.
We then show that there always exists an optimal solution in which at most two choices have positive values.
@@
$(1)\quad w_{ij^{\prime}}\geq w_{ij}\qquad(2)\quad\mu_{ij^{\prime}}\leq\mu_{ij}$
@@
First, if $j^{\prime}$ single-dominates $j$, define a new solution by setting $x^{*}_{ij^{\prime}}=x_{ij^{\prime}}+x_{ij},\ x^{*}_{ij}=0.$
@@
This again yields a feasible solution whose objective value is at least as large as that of $x$ (see fig. 2 for illustration, and section D.2 for the full proof). Therefore, in either case, there exists an optimal solution in which $x_{ij}=0$.
@@
There exists
an optimal solution $x$ to $\Phi(t)$ in which, for every item $i$, at most two choices satisfy $x_{ij}>0$.
@@
Let $x$ be an optimal solution in which no choice is dominated; such a solution exists by Lemma 3.4. Without loss of generality, we may further assume that no additional choices remain and that for every $i\in[n]$ and $j\in[m-1]$, $w_{i,j}<w_{i,j+1}$. Lemma 3.4 then also yields $\mu_{i,j}<\mu_{i,j+1}$.
@@
Suppose for some item $i$ there are strictly more than two choices that satisfy $x_{ij}>0$.
Define $\tilde{\mu}=\sum_{j}x_{ij}\mu_{ij}$. Let $j^{\prime}=\arg\max_{j}\bigl\{\mu_{ij}\leq\tilde{\mu}\bigr\}$ and $j^{\prime\prime}=\arg\min_{j}\bigl\{\mu_{ij}\geq\tilde{\mu}\bigr\}$. By the definition of $\tilde{\mu}$ and by our assumption that at least two choices satisfy $x_{ij}>0$, such indices $j^{\prime}$ and $j^{\prime\prime}$ always exist. Set $\lambda=\frac{\mu_{ij^{\prime\prime}}-\tilde{\mu}}{\mu_{ij^{\prime\prime}}-\mu_{ij^{\prime}}}$. Consider the solution defined by $x^{*}_{ij^{\prime}}=\lambda$, $x^{*}_{ij^{\prime\prime}}=1-\lambda$, and $x^{*}_{ij}=0$ for all $j\notin\{j^{\prime}j^{\prime\prime}\}$. Note that this construction satisfies $\sum_{j}x^{*}_{ij}\,\mu_{ij}=\tilde{\mu}$. Therefore, it yields a feasible solution. Since the resulting point lies on the boundary of the convex set, its objective value is at least as large as that of the original solution $x$, (see fig. 3 for illustration).
@@
By Lemmas 3.4, 3.5, it follows that there exists an optimal solution $x$ where at most two choices have positive values.
@@
Figure 4: For an optimal solution $x$, items $i_{1}$ and $i_{2}$ each have two choices with nonzero values. The solution for each item is a convex combination of the two choices, corresponding to the red dots in the graphs $\Gamma_{i_{1}}$ and $\Gamma_{i_{2}}$, which lie on the segments connecting the two choices. As the slope of the first item is larger, the modified solution indicated by the dashed arrows illustrates one possible update that maintains feasibility and increases the objective value.
For a set of item choices $\mathcal{S}\subset[n]\times[m]$, denote $size(\mathcal{S})=\sum_{(i,j)\in\mathcal{S}}s_{i}^{j}$ and $\mu(S)=\sum_{(i,j)\in\mathcal{S}}\mu_{ij}$. The following Lemmas are direct extensions of key lemmas from Dean et al. [10], originally established for the classic Stochastic Knapsack problem with a capacity of 1. We adapt these lemmas to our knapsack variant with multi-choice and costs, and for a general capacity $t$. In this section, we only need the case where $t=1$; however, in appendix G we use the generalized versions. The proofs are presented in Appendix D for completeness.
@@
For any set of item choices $\mathcal{S}$, $\Pr[size(\mathcal{S})<t]\geq 1-\frac{1}{t}\mu(\mathcal{S}).$
@@
For any adaptive policy for the MSKC problem with capacity $t$, let $\mathcal{S}$ be the random set of item choices that the policy inserts, then $\operatorname{\mathbb{E}}[\mu(\mathcal{S})]\leq 2t$.
@@
For any constant $c\geq 2$ and capacity $t$, $\mathrm{ADAPT}\leq\Phi(ct)\leq c\cdot\Phi(t),$ where $\mathrm{ADAPT}$ is the optimal adaptive policy for capacity $t$.
We now present our algorithm, NA-SKC (Non-Adaptive Stochastic Knapsack with Costs):
@@
Solve linear program $\Phi(1)$ to obtain optimal solution $x$ with the structure from Lemma 3.3 Let $\mathcal{C}=\{(i,j)\mid x_{ij}>0\}$ be the candidate item-choice pairs
@@
   Remove $(i^{*},j)$ and $(i^{*},j^{\prime})$ from $\mathcal{C}$ Create virtual item $(i^{*},j^{*})$ with:
@@
   Add $(i^{*},j^{*})$ to $\mathcal{C}$ When attempting to insert $(i^{*},j^{*})$, choose $j$ with probability $x_{i^{*},j}$ and $j^{\prime}$ with probability $x_{i^{*},j^{\prime}}$